# thanks: Corresponding author

### v1

@@ Abstract
We present a detailed analysis of the proton electromagnetic form factors (FFs) using exclusive photon leptoproduction (EP) data in kinematic regions where the Bethe-Heitler (BH) contribution dominates the deeply virtual Compton scattering (DVCS) cross section. By exploiting the strong sensitivity of the BH amplitude to the Dirac and Pauli FFs, we determine $F_{1}(t)$, $F_{2}(t)$ and the corresponding Sachs FFs from several fitting scenarios based on dipole and $P$-pole parametrizations, and evaluate the charge and magnetic radii of the proton. We show that EP measurements in the range $0.11<|t|<0.45$ GeV2 provide meaningful constraints on $F_{1}(t)$, while offering limited sensitivity to $F_{2}(t)$. Our results systematically favor smaller charge radius values compared with those from traditional elastic electron-proton scattering measurements, consistent with recent high-precision PRad measurements. This study demonstrates that EP measurements, especially when covering smaller values of $|t|$, can serve as a complementary and powerful tool for determining the proton electromagnetic structure and may offer additional insight into the long-standing charge-radius puzzle. The methodology developed here establishes a framework for future combined analyses of EP and elastic electron-proton scattering data which enables a unified determination of the nucleon FFs.
@@ I Introduction
It has been well stabilised now that deeply virtual Compton scattering (DVCS), an exclusive process where a lepton (e.g., electron) scatters off a hadron (e.g., proton) by exchanging a virtual photon and a real photon is produced in the final state, remaining the target intact, is one of the most powerful tools for probing the internal structure of hadrons Ji:1996nm; Belitsky:2001ns. In contrast to inclusive deep inelastic scattering (DIS), which provides only information on longitudinal momentum distributions of partons through the parton distribution functions (PDFs) Alekhin:2002fv, DVCS provides access to generalized parton distributions (GPDs) Ji:1996ek; Ji:1996nm; Radyushkin:1996nd; Radyushkin:1997ki; Collins:1998be; Belitsky:2001ns; Guidal:2013rya; Kumericki:2016ehc; Aschenauer:2025cdq; Guo:2025muf; Xie:2023xkz. GPDs are more comprehensive objects compared to PDFs and represent the correlations between the spatial and momentum degrees of freedom inside the nucleon Diehl:2003ny; Ji:2004gf; Belitsky:2005qn; Boffi:2007yc; Diehl:2015uka; Kriesten:2020wcx; Mezrag:2022pqk; Lorce:2025aqp. Through these distributions, DVCS can provide insights into fundamental aspects of quantum chromodynamics (QCD), such as the three-dimensional imaging of the nucleon through GPDs Burkardt:2002hr; Kaur:2023lun; Cichy:2024afd, the decomposition of the nucleon spin among its quark and gluon constituents Aidala:2012mv; Ji:2020ena, the connection between various nucleon form factors (FFs) and partonic correlations Diehl:2003ny, and the mechanical properties of the nucleon Polyakov:2018zvc; Burkert:2018bqq; Kumericki:2019ddg; Martinez-Fernandez:2025jvk.
@@
The theoretical framework of QCD factorization ensures that the DVCS amplitude can be expressed in terms of perturbatively calculable coefficient functions and universal GPDs, which enter through the so-called Compton form factors (CFFs) Collins:1998be; Guo:2023ahv; CaleroDiaz:2025luc. Experimentally, DVCS has been investigated extensively at HERA, JLab, COMPASS, and HERMES, leading to a wealth of precise measurements of cross sections and related observables such as beam-spin, beam-charge, and target-spin asymmetries Kumericki:2016ehc. These observables are highly sensitive to the underlying GPDs. Therefore, the precise measurements of DVCS observables can be used in global analyses in order to extract the functional form of GPDs Kumericki:2016ehc; Kriesten:2019jep; Guo:2023ahv. Moreover, DVCS is often studied in conjunction with the Bethe-Heitler (BH) process, in which the real photon is emitted from the lepton rather than the nucleon. The interference between the DVCS and BH amplitudes plays a central role in unravelling the nucleon structure Kriesten:2019jep.
@@
GPDs are directly linked, through their Mellin moments, to a wide class of hadronic FFs, including electromagnetic Guidal:2004nd; Diehl:2013xca; Hashamipour:2022noy, axial Hashamipour:2019pgy; Irani:2023lol; Bhattacharya:2024wtg, gravitational Pagels:1966zza; Polyakov:2018zvc; Goharipour:2025lep, and transition FFs Goharipour:2024atx. Among these, the electromagnetic FFs have a central role in hadron structure studies. They are commonly expressed either in terms of the electric and magnetic Sachs FFs, $G_{E}(t)$ and $G_{M}(t)$, or equivalently in terms of the Dirac and Pauli FFs, $F_{1}(t)$ and $F_{2}(t)$, where $t$ is the squared four-momentum transfer. These quantities encode the spatial distributions of electric charge and magnetic moment within the nucleon. In particular, precise measurements of these FFs allow one to determine the nucleon charge and magnetic radii, $r_{E}$ and $r_{M}$, which are directly related to the slopes of the FFs at $t=0$ Karr:2020wgh; Gao:2021sml; Xiong:2023zih. Such measurements have attracted great interest in recent years, due to the so-called proton charge radius puzzle Lorenz:2012tm; Pohl:2013yb; Bernauer:2014cwa; Carlson:2015jba; Hill:2017wzi; Peset:2021iul; Lin:2023fhr; Lumpay:2025btu, which refers to the discrepancy observed in measurements of the proton’s charge radius when using different methods. Beyond their static interpretation, electromagnetic FFs can also be considered as important benchmarks for testing various theoretical and phenomenological approaches to QCD, such as lattice simulations Park:2021ypf, holographic QCD Xu:2021wwj, and phenomenological GPD models Goharipour:2024mbk (see Ref. Goharipour:2025yxm and references therein for more information). Consequently, the extraction of $F_{1}$ and $F_{2}$ from experimental observables remains an important mission in hadronic physics.
@@
To extract the electromagnetic FFs from the experimental data, the measurements of the elastic electron-nucleon scattering cross section is usually used Gao:2003ag; Qattan:2004ht; Perdrisat:2006hj; Crawford:2006rz; Arrington:2007ux; A1:2010nsl; A1:2013fsc; Ye:2017gyb; Xiong:2019umf; Christy:2021snt; Williams:2025fiv. In the present study, we aim to examine the possible extraction of these FFs from the measurements of the exclusive photon leptoproduction (EP) process which comprises the DVCS and BH amplitudes as well as their interference. To be more precise, since the BH contribution to the EP process is directly sensitive to the electromagnetic FFs, a novel idea to extract them is to analyse EP data in kinematic regions where the BH term is strongly dominant over the other contributions. To this aim, we first examine this idea by calculating different contributions in the EP cross section separately for a nominal energy scale $Q^{2}=1$ GeV2 and searching for the kinematic regions where the BH contribution is dominant. Then, we extract Dirac and Pauli FFs, $F_{1}(t)$ and $F_{2}(t)$, by analyzing the JLab data considering only data points for which the BH contribution is dominant.
@@ II EP cross section and BH contribution
One of the most insightful approaches to access GPDs is through exclusive processes Ji:1998pc; Goeke:2001tz; Kroll:2012sm; Duplancic:2022ffo; Qiu:2022pla. Exclusive processes are characterized by the requirement that the hadron remains intact in the final state. Among the most important of these processes are DVCS Ji:1996ek; Ji:1996nm; Radyushkin:1996nd; Radyushkin:1997ki; Collins:1998be; Belitsky:2001ns; Guidal:2013rya; Kumericki:2016ehc; Aschenauer:2025cdq; Guo:2025muf; Xie:2023xkz and deeply virtual meson production (DVMP) Collins:1996fb; Goloskokov:2005sd; Goloskokov:2007nt; Muller:2013jur; Favart:2015umi; Cuic:2023mki; Boussarie:2024pax; Hatta:2025vhs. In both cases, the target remains intact, while an additional particle is produced: a real photon in DVCS and a meson in DVMP. In the present study, we focus on EP, which is described by the following process:
@@
Figure 1: EP process shown in terms of laboratory-frame kinematic variables. Figure taken from Ref. Belitsky:2001ns.
The first computation of the DVCS cross section in terms of leading-twist CFFs was carried out by Ji Ji:1996nm. Subsequently, Belitsky, Müller, and Kirchner (BKM) Belitsky:2001ns made a substantial contribution by extending the analysis beyond the leading-order treatment of GPDs. They provided explicit formulas for DVCS cross section in terms of laboratory-frame kinematic variables, as illustrated in Fig. 1. In their convention, the azimuthal angle $\phi_{N}\equiv\phi$ denotes the orientation of the recoiling hadron. A convenient reference frame is chosen such that the virtual photon four-momentum $q_{1}$ has no transverse components. The $z$-axis is fixed along the negative direction of $q_{1}$, while the incoming lepton has a positive $x$-component. Explicitly, the relevant four-momenta are defined as
@@
$P_{2}$ $=\left(E_{2},\,|P_{2}|\cos\phi_{N}\sin\theta_{N},\,|P_{2}|\sin\phi_{N}\sin\theta_{N},\,|P_{2}|\cos\theta_{N}\right)\,.$ (2)
In this frame, $\theta_{\ell}$ denotes the polar angle of the incoming lepton with respect to the negative $z$-axis (the virtual photon direction), while $\theta_{N}$ represents the polar angle of the recoiling proton relative to the same axis. The azimuthal angle of the scattered lepton is fixed to $\phi_{\ell}=0$, while the angle between the lepton plane and the recoiling proton momentum is given by $\phi_{N}$. Here $E$ is the energy of the incoming lepton, $E_{2}$ the energy of the recoiling proton, and $M$ the proton mass. Finally, the longitudinal polarization vector is taken as $S_{LP}=\left(0,\,0,\,0,\,\Lambda\right)$.
In addition to DVCS, there is another process with identical initial and final states, which cannot be experimentally distinguished, known as the BH process Belitsky:2001ns. Figure 2 illustrates the DVCS diagram (left) together with the two BH diagrams (center and right). In the BH case, the incoming electron is elastically scattered by the nucleon, while the real photon is emitted from the lepton line rather than from the struck quark, either after (center) or before (right) the lepton scatters off the nucleon. The four-momenta are assigned as $q\equiv q_{1}$ for the virtual photon, $q^{\prime}\equiv q_{2}$ for the real photon, and $\Delta=p^{\prime}-p=q-q^{\prime}$ with $t=\Delta^{2}$, where $p\equiv P_{1}$ and $p^{\prime}\equiv P_{2}$. Unlike DVCS, the BH process is entirely determined by the elastic nucleon FFs, $F_{1}(t)$ and $F_{2}(t)$ Belitsky:2001ns.
Figure 2: Diagrammatic representation of DVCS and BH processes. In DVCS (left), the real photon is emitted by the struck quark inside the proton, whereas in BH it is emitted from the lepton line, either after (center) or before (right) scattering off the proton. Figure taken from Ref. Adams:2024pxw.
@@
The four-fold differential cross section of EP can be expressed as Belitsky:2010jw; Benali:2020vma:
@@
The BH amplitude, calculated within the framework of pure QED, can be expressed in terms of the nucleon elastic FFs. The squared amplitude of BH is given by the following expression Belitsky:2001ns:
@@
where $P_{1}(\phi)$ and $P_{2}(\phi)$ denote the lepton propagators corresponding to the two BH diagrams. The complete expressions for these propagators and the harmonic coefficients $c_{i}^{\text{BH}}$ and $s_{1}^{\text{BH}}$ are provided in Ref. Belitsky:2001ns. For an unpolarized target, the sine coefficient vanishes, i.e., $s_{1}^{\text{BH}}=0$. The harmonic coefficients are functions of the proton electromagnetic FFs $F_{1}(t)$ and $F_{2}(t)$. This dependence provides a unique opportunity: in kinematic regions where the BH contribution dominates over the pure DVCS and interference terms, the measured cross section becomes directly sensitive to the proton’s electromagnetic FFs. This observation underlies the central motivation of the present study.
The explicit dependence of the BH cross section on the electromagnetic FFs becomes particularly transparent in an alternative formalism proposed in Refs. Kriesten:2019jep, where the BH contribution is cast in a Rosenbluth-type representation, analogous to that used in elastic lepton-nucleon scattering. In this framework, the unpolarized BH cross section naturally separates into terms proportional to the Dirac and Pauli FFs, that makes the individual roles of $F_{1}$ and $F_{2}$ explicit. This decomposition provides direct sensitivity to the underlying elastic structure of the nucleon even within the deeply virtual kinematic regime. The unpolarized BH contribution to the four-fold differential cross section can be written as Kriesten:2019jep; Adams:2024pxw:
@@
with $\gamma=4M^{2}x_{B}^{2}/Q^{2}$ and $s=(k+p)^{2}$ denoting the electron-proton center-of-mass energy squared. Note that the Dirac and Pauli FFs are related to the Sachs FFs as $G_{E}=F_{1}-\tau F_{2}$ and $G_{M}=F_{1}+F_{2}$. The coefficients $A_{\text{BH}}$ and $B_{\text{BH}}$ are dimensionless kinematic functions of $(s,Q^{2},x_{B},t,\phi)$. Their explicit expressions were derived in Ref. Kriesten:2019jep, where they are presented in a fully covariant form. Equation (7) thus demonstrates how the BH term can be organized in a Rosenbluth-like framework, providing a transparent and model-independent link between the measurable cross section and the elastic nucleon FFs.

### v2

@@ Abstract
We present a detailed analysis of the proton electromagnetic form factors (FFs) using exclusive photon leptoproduction (EP) data in kinematic regions where the Bethe–Heitler (BH) contribution dominates the deeply virtual Compton scattering (DVCS) cross section. By exploiting the sensitivity of the BH amplitude to the Dirac and Pauli FFs, we extract $F_{1}(t)$, $F_{2}(t)$, and the corresponding Sachs FFs within several fitting scenarios based on dipole and $P$-pole parametrizations, and evaluate the charge and magnetic radii of the proton. In this fitting scenario, we show that EP measurements in the range $0.11<|t|<0.45~\mathrm{GeV}^{2}$ can provide constraints on $F_{1}(t)$, while offering limited sensitivity to $F_{2}(t)$. The extracted charge radius values tend to be smaller than those obtained from traditional elastic electron–proton scattering measurements and are consistent, within uncertainties, with recent high-precision PRad results. These findings indicate that EP measurements, especially when covering smaller values of $|t|$, can serve as a complementary tool for determining the proton electromagnetic structure and may contribute to ongoing efforts to better understand the proton charge radius. The methodology developed here provides a framework for future combined analyses of EP and elastic electron-proton scattering data which enables a unified determination of the nucleon FFs.
@@ I Introduction
It has been well stabilized now that deeply virtual Compton scattering (DVCS), an exclusive process where a lepton (e.g., electron) scatters off a hadron (e.g., proton) by exchanging a virtual photon and a real photon is produced in the final state, leaving the target intact, is one of the most powerful tools for probing the internal structure of hadrons Ji (1997a); Belitsky et al. (2002). In contrast to inclusive deep inelastic scattering (DIS), which provides only information on longitudinal momentum distributions of partons through the parton distribution functions (PDFs) Alekhin (2003), DVCS provides access to generalized parton distributions (GPDs) Ji (1997b, a); Radyushkin (1996, 1997); Collins and Freund (1999); Belitsky et al. (2002); Guidal et al. (2013); Kumericki et al. (2016); Aschenauer and others (2025); Guo et al. (2025); Xie et al. (2023); Braun et al. (2025). GPDs are more comprehensive objects compared to PDFs and represent the correlations between the spatial and momentum degrees of freedom inside the nucleon Diehl (2003); Ji (2004); Belitsky and Radyushkin (2005); Boffi and Pasquini (2007); Diehl (2016); Kriesten and Liuti (2022); Mezrag (2022); Lorcé et al. (2025); Boër and others (2025). Through these distributions, DVCS can provide insights into fundamental aspects of quantum chromodynamics (QCD), such as the three-dimensional imaging of the nucleon through GPDs Burkardt (2003); Kaur et al. (2024); Cichy et al. (2024), the decomposition of the nucleon spin among its quark and gluon constituents Aidala et al. (2013); Ji et al. (2021), the connection between various nucleon form factors (FFs) and partonic correlations Diehl (2003), and the mechanical properties of the nucleon Polyakov and Schweitzer (2018); Burkert et al. (2018); Kumerički (2019); Martínez-Fernández et al. (2025).
@@
The theoretical framework of QCD factorization ensures that the DVCS amplitude can be expressed in terms of perturbatively calculable coefficient functions and universal GPDs, which enter through the so-called Compton form factors (CFFs) Collins and Freund (1999); Guo et al. (2023); Calero Diaz and Keller (2025). Experimentally, DVCS has been investigated extensively at HERA, JLab, COMPASS, and HERMES, leading to a wealth of precise measurements of cross sections and related observables such as beam-spin, beam-charge, and target-spin asymmetries (see Ref. Kumericki et al. (2016) and references therein). These observables are highly sensitive to the underlying GPDs. Therefore, the precise measurements of DVCS observables can be used in global analyses in order to extract the functional form of GPDs Kumericki et al. (2016); Kriesten et al. (2020); Guo et al. (2023). Moreover, measurements of DVCS necessarily include the Bethe-Heitler (BH) process, in which the real photon is emitted from the lepton rather than the nucleon, and which often dominates the cross section in the experimentally accessible kinematic regions. The interference between the DVCS and BH amplitudes plays a central role in unraveling the nucleon structure Kriesten et al. (2020).
@@
GPDs are directly linked, through their Mellin moments, to a wide class of hadronic FFs, including electromagnetic Guidal et al. (2005); Diehl and Kroll (2013); Hashamipour et al. (2023), axial Hashamipour et al. (2019); Irani et al. (2023); Bhattacharya et al. (2025), gravitational Pagels (1966); Polyakov and Schweitzer (2018); Goharipour et al. (2025a), and transition FFs Goharipour et al. (2024). Among these, the electromagnetic FFs have a central role in hadron structure studies. They are commonly expressed either in terms of the electric and magnetic Sachs form factors, $G_{E}(t)$ and $G_{M}(t)$, or equivalently in terms of the Dirac and Pauli form factors, $F_{1}(t)$ and $F_{2}(t)$, where $t$ is the Mandelstam variable corresponding to the squared four-momentum transfer to the nucleon. In the kinematic region considered in this work, $t$ is negative. These quantities encode the spatial distributions of electric charge and magnetic moment within the nucleon. In particular, precise measurements of these FFs allow one to determine the nucleon charge and magnetic radii, $r_{E}$ and $r_{M}$, which are directly related to the slopes of the FFs at $t=0$ Karr et al. (2020); Gao and Vanderhaeghen (2022); Xiong and Peng (2023). Such measurements have attracted great interest in recent years, due to the so-called proton charge radius puzzle Lorenz et al. (2012); Pohl et al. (2013); Bernauer and Pohl (2014); Carlson (2015); Hill (2017); Peset et al. (2021); Lin et al. (2024); Lumpay et al. (2025), which refers to the discrepancy observed in measurements of the proton’s charge radius when using different methods. Beyond their static interpretation, electromagnetic FFs can also be considered as important benchmarks for testing various theoretical and phenomenological approaches to QCD, such as lattice simulations Park et al. (2022), holographic QCD Xu et al. (2021), and phenomenological GPD models Goharipour et al. (2025c) (see Ref. Goharipour et al. (2025b) and references therein for more information). Consequently, the extraction of $F_{1}$ and $F_{2}$ from experimental observables remains an important mission in hadronic physics.
@@
To extract the electromagnetic FFs from the experimental data, the measurements of the elastic electron-nucleon scattering cross section is usually used Gao (2003); Qattan and others (2005); Perdrisat et al. (2007); Crawford and others (2007); Arrington et al. (2007); Bernauer and others (2010, 2014); Ye et al. (2018); Xiong and others (2019); Christy and others (2022); Williams et al. (2025); Vaziri et al. (2025). In the present study, we aim to examine the possible extraction of these FFs from the measurements of the exclusive photon leptoproduction (EP) process which comprises the DVCS and BH amplitudes as well as their interference. To be more precise, since the BH contribution to the EP process is directly sensitive to the electromagnetic FFs, a novel idea to extract them is to analyse EP data in kinematic regions where the BH term is strongly dominant over the other contributions. To this aim, we first examine this idea by calculating different contributions in the EP cross section separately for a nominal energy scale $Q^{2}=1$ GeV2 and searching for the kinematic regions where the BH contribution is dominant. Then, we extract Dirac and Pauli FFs, $F_{1}(t)$ and $F_{2}(t)$, by analyzing the JLab data considering only data points for which the BH contribution is dominant.
@@ II EP cross section and BH contribution
One of the most insightful approaches to access GPDs is through exclusive processes Ji (1998); Goeke et al. (2001); Kroll et al. (2013); Duplančić et al. (2023); Qiu and Yu (2023). Exclusive processes are characterized by the requirement that the hadron remains intact in the final state. Among the most important of these processes are DVCS Ji (1997b, a); Radyushkin (1996, 1997); Collins and Freund (1999); Belitsky et al. (2002); Guidal et al. (2013); Kumericki et al. (2016); Aschenauer and others (2025); Guo et al. (2025); Xie et al. (2023); Braun et al. (2025) and deeply virtual meson production (DVMP) Collins et al. (1997); Goloskokov and Kroll (2005, 2008); Müller et al. (2014); Favart et al. (2016); Čuić et al. (2023); Boussarie et al. (2025); Hatta et al. (2025). In both cases, the target remains intact, while an additional particle is produced: a real photon in DVCS and a meson in DVMP. In the present study, we focus on EP, which is the experimentally accessible process underlying both DVCS and BH mechanisms. The EP process is described by
@@
Figure 1: EP process shown in terms of laboratory-frame kinematic variables. Figure taken from Ref. Belitsky et al. (2002).
The first computation of the DVCS cross section in terms of leading-twist CFFs was carried out by Ji Ji (1997a). Subsequently, Belitsky, Müller, and Kirchner (BKM) Belitsky et al. (2002) made a substantial contribution by extending the analysis beyond the leading-order treatment of GPDs. They provided explicit formulas for DVCS cross section in terms of laboratory-frame kinematic variables, as illustrated in Fig. 1. In their convention, the azimuthal angle $\phi$ denotes the orientation of the recoiling hadron. Henceforth, we adopt the notation $\phi$ to denote the azimuthal angle of the recoiling hadron. Other relevant kinematic quantities, such as the polar and azimuthal angles of the lepton, which are not explicitly shown in Fig. 1, are defined below as needed. A convenient reference frame is chosen such that the virtual photon four-momentum $q_{1}$ has no transverse components. The $z$-axis is fixed along the negative direction of $q_{1}$, while the incoming lepton has a positive $x$-component. Explicitly, the relevant four-momenta are defined as
@@
$P_{2}$ $=\left(E_{2},\,|P_{2}|\cos\phi\sin\theta_{N},\,|P_{2}|\sin\phi\sin\theta_{N},\,|P_{2}|\cos\theta_{N}\right)\,.$ (2)
In this frame, $\theta_{\ell}$ denotes the polar angle of the incoming lepton with respect to the negative $z$-axis (the virtual photon direction), while $\theta_{N}$ represents the polar angle of the recoiling proton relative to the same axis. The azimuthal angle of the scattered lepton is fixed to $\phi_{\ell}=0$, while, as defined in Fig. 1, the angle between the lepton plane and the recoiling proton momentum is given by $\phi$. Here $E$ is the energy of the incoming lepton, $E_{2}$ the energy of the recoiling proton, and $M$ the proton mass. Finally, the longitudinal polarization vector is taken as $S_{LP}=\left(0,\,0,\,0,\,\Lambda\right)$. $\varphi$ represents the angle between this vector and the hadronic plane.
In addition to DVCS, there is another process with identical initial and final states, which cannot be experimentally distinguished, known as the BH process Belitsky et al. (2002). Figure 2 illustrates the DVCS diagram (left) together with the two BH diagrams (center and right). In the BH case, the incoming electron is elastically scattered by the nucleon, while the real photon is emitted from the lepton line rather than from the struck quark, either after (center) or before (right) the lepton scatters off the nucleon. The four-momenta are assigned as $q\equiv q_{1}$ for the virtual photon, $q^{\prime}\equiv q_{2}$ for the real photon, and $\Delta=p^{\prime}-p=q-q^{\prime}$ with $t=\Delta^{2}$, where $p\equiv P_{1}$ and $p^{\prime}\equiv P_{2}$. Unlike DVCS, the BH process is entirely determined by the elastic nucleon FFs, $F_{1}(t)$ and $F_{2}(t)$ Belitsky et al. (2002).
Figure 2: Diagrammatic representation of DVCS and BH processes. In DVCS (left), the real photon is emitted by the struck quark inside the proton, whereas in BH it is emitted from the lepton line, either after (center) or before (right) scattering off the proton. Figure taken from Ref. Adams and others (2024).
@@
The four-fold differential cross section of EP can be expressed as Belitsky and Mueller (2010); Benali and others (2020):
@@
The BH amplitude, calculated within the framework of pure QED, can be expressed in terms of the nucleon elastic FFs. The squared amplitude of BH is given by the following expression Belitsky et al. (2002):
@@
where $P_{1}(\phi)$ and $P_{2}(\phi)$ denote the lepton propagators corresponding to the two BH diagrams. The complete expressions for these propagators and the harmonic coefficients $c_{i}^{\text{BH}}$ and $s_{1}^{\text{BH}}$ are provided in Ref. Belitsky et al. (2002). For an unpolarized target, the sine coefficient vanishes, i.e., $s_{1}^{\text{BH}}=0$. The harmonic coefficients are functions of the proton electromagnetic FFs $F_{1}(t)$ and $F_{2}(t)$. This dependence provides a unique opportunity: in kinematic regions where the BH contribution dominates over the pure DVCS and interference terms, the measured cross section becomes directly sensitive to the proton’s electromagnetic FFs. This observation underlies the central motivation of the present study.
The explicit dependence of the BH cross section on the electromagnetic FFs becomes particularly transparent in an alternative formalism proposed in Refs. Kriesten et al. (2020), where the BH contribution is cast in a Rosenbluth-type representation, analogous to that used in elastic lepton-nucleon scattering. In this framework, the unpolarized BH cross section naturally separates into terms proportional to the Dirac and Pauli FFs, that makes the individual roles of $F_{1}$ and $F_{2}$ explicit. This decomposition provides direct sensitivity to the underlying elastic structure of the nucleon even within the deeply virtual kinematic regime. The unpolarized BH contribution to the four-fold differential cross section can be written as Kriesten et al. (2020); Adams and others (2024):
@@
with $\gamma=4M^{2}x_{B}^{2}/Q^{2}$ and $s=(k+p)^{2}$ denoting the electron-proton center-of-mass energy squared. Note that the Dirac and Pauli FFs are related to the Sachs FFs as $G_{E}=F_{1}-\tau F_{2}$ and $G_{M}=F_{1}+F_{2}$. The coefficients $A_{\text{BH}}$ and $B_{\text{BH}}$ are dimensionless kinematic functions of $(s,Q^{2},x_{B},t,\phi)$. Their explicit expressions were derived in Ref. Kriesten et al. (2020), where they are presented in a fully covariant form. Equation (7) thus demonstrates how the BH term can be organized in a Rosenbluth-like framework, providing a transparent and model-independent link between the measurable cross section and the elastic nucleon FFs.